# Earth Observation Foundation Model PhilEO: Pretraining on the MajorTOM and FastTOM Datasets

### v2

@@ I Introduction
EO FMs can be trained, for example, using MAE [1], [56], [57], contrastive learning [2], or Joint Embedding Predictive Architecture (JEPA) [39]. MAE refers to masking (e.g., a large percentage of the image such as $75\%$), using the reconstruction loss, and the MSE [23], [26]. Also, the ViT [9] (or a variant of the ViT like the Swin Transformer [12]) is usually used. Here, we also note that the Spectral Angle Mapper (SAM), which takes into account the spectral signatures of the observations, can also be used in addition to the MSE, or even on its own as the loss function that is minimized during training. In addition, contrastive learning refers to performing data augmentation and bringing closer together in the latent feature space the data-augmented samples, repelling the dissimilar samples.
Our scaling-up study. In this paper, we present the scaling-up of our recently proposed model PhilEO on the unlabeled 23TB dataset MajorTOM [15], which covers the vast majority of the Earth’s surface, as well as on the specialized subset FastTOM 2TB that does not include oceans and ice. FastTOM is a task-specific dataset for land only. In the literature, there exist FMs only for oceans or ice, for example HydroFM [25], but we are interested in Earth Observation and not only in Land Observation or Ocean Observation (or Ice Observation). To this end, in this work, we develop and study various models with different numbers of parameters, and also with different architectures. We evaluate and examine the FLoating-point OPerations (FLOPs) that are needed by the models. We fine-tune the models on the PhilEO Bench [3] for road density estimation, building density pixel-wise regression, and land cover semantic segmentation, and we evaluate their performance.
@@
  • We scale the pretraining from 0.5TB to 23TB, which is a big scale-up, as well as to 2TB, where the 23TB are general-purpose EO data (i.e. MajorTOM [^1] http://github.com/ESA-PhiLab/Major-TOM and also in http://huggingface.co/Major-TOM), which include land, oceans, and ice, while the 2TB are a specialized subset that does not include oceans and ice. We use different number of model parameters and architectures [^2] We release the code of our model in: http://github.com/ESA-PhiLab/PhilEO-MajorTOM.. We show that for the task of road density regression, for the $n$-shot $n=500$, larger pretraining data reduces the final downstream task’s Root Mean Squared Error (RMSE), resulting in a percentage improvement $12.0\%$.
  • In addition to pretraining dataset scaling, we also study the scaling-up of the model parameters and the architecture, comparing CNN- and ViT-based models, as well as Mamba. We evaluate the effectiveness of the scaling-up on the PhilEO Bench, covering both road and building density estimation, as well as land cover mapping.
  • We develop and compare different architectures for the EO Foundation Models in order to select which one to scale-up. More specifically, for road density regression, we show that the ViT UPerNet model outperforms the ViT with a convolutional decoder in RMSE, with a percentage improvement of $34.2\%$ for the $n$-shot $n=100$. On the same task, the U-Net CNN Geo-Aware model achieves the same performance as the ViT UPerNet.
@@ II Background
MajorTOM and FastTOM datasets. The MajorTOM dataset [15] comprises in total approximately 60TB of unlabeled global data, covering most of the Earth, including oceans and ice. In this work, we focus on the MajorTOM Core-S2L2A subset, 23TB. In addition, FastTOM is a 2TB specialized subset of MajorTOM Core-S2L2A, containing only land and excluding oceans and ice, making it more task-specific for terrestrial downstream tasks.
@@ III-B PhilEO ViT UPerNet
ViT-UPerNet architectures have been shown to achieve state-of-the-art performance in various computer vision tasks [11]-[14], including segmentation and classification. For models that use a ViT backbone, we extract intermediate feature maps, $\{C_{2},C_{3},C_{4},C_{5}\}$ in Fig. 1, across different layers. These feature maps are then appropriately resized (i.e. upsampled or downsampled) to align spatial dimensions.
The highest-level feature map $C_{5}$ is processed by the PPM, which applies pooling operations at multiple scales to capture contextual information at different resolutions. The outputs of the PPM and the FPN are then fused through a series of 1×1 convolutions and element-wise additions, producing the final set of multi-scale feature maps, $\{P_{2},P_{3},P_{4},P_{5}\}$ in Fig. 1. These aggregated features are then used for downstream pixel-wise prediction tasks, including regression and segmentation.
@@ III-C Mamba 2D Models
ViT models have significantly less image-inductive bias compared to CNN U-Net models. This is because ViT attention correlates everything with everything else. Moreover, in self-attention, image token order does not matter [9], [30]. In addition, Mamba models have more image-inductive bias compared to ViT models, as in Mamba S6 SSM models, the image token order is important and matters.
@@ IV Related Work
Recent studies on GFMs, summarized in Table I, highlight the increasing potential of learning robust feature representations from large-scale EO data, later used for fine-tuning on various downstream tasks. For a recent study on the scope of geo-spatial FMs, we refer the reader to [60]. However, most (but certainly not all) existing studies have been limited by relatively modest dataset sizes and have not explored training GFMs at the scale of more than 20TB, and this is important. This motivates our work, where we develop and pretrain the PhilEO MajorTOM model on the full 23TB MajorTOM Core-S2L2A dataset. Furthermore, during the ESA-NASA Workshop on AI Foundation Models for Earth Observation [61] (in May 2025), it was highlighted that the performance of FMs is not uniform across tasks; moreover, in certain cases, conventional models surpass state-of-the-art EO FMs. This observation prompts a critical inquiry: should the field pursue a single universal model, develop application-specific models, or adopt a Mixture of Experts (MoE) strategy?
@@ V Evaluation and Results
In this section, we present the evaluation and results of the models when we scale-up EO FMs in terms of larger pretraining dataset, i.e. the MajorTOM Core-S2L2A 23TB dataset, larger model, and architecture. According to the results, evaluation, and experiments, EO FMs pretrained on general-purpose data, when scaled-up appropriately, can eventually outperform models pretrained on specialized data. We are interested in Earth Observation, and not only in Land Observation or Ocean Observation (or Ice Observation). In the literature, there exist FMs only for oceans or ice, for example HydroFM [25]. When we pretrain on the entire MajorTOM 23TB dataset, we do not use any prior information, while when we pretrain on specialized datasets that have only land and exclude oceans, ice, and deserts, i.e. the FastTOM 2TB dataset, then we use prior information. The downstream tasks we are interested in are on land, i.e. PhilEO Bench [3]: road density estimation, building density regression, and land cover mapping [^4] The PhilEO Bench downstream tasks labeled dataset (0.40.4TB) was presented in [3], and we also note that a small and more manageable subset can be found in http://huggingface.co/datasets/JenteBosmans/tiny_phileo.. We focus on land downstream tasks and not on oceans or ice (or deserts), and this is why specialized datasets have prior knowledge.
Figure 4: Road density estimation: Evaluation in RMSE at different $n$-shots, where the best performing model is PhilEO Geo-Aware U-Net. The PhilEO ViT UPerNet black line is below the Geo-Aware U-Net red line. PhilEO ViT UPerNet outperforms ViT CNN and ViT CNN Grouped Channels (GC) [3].
@@ V-B Scaling-up: Evaluation of PhilEO MajorTOM 23TB GFM
For the evaluation of the models, continuing the discussion in the preceding paragraph, we count the FLOPs that are needed by the different models. The lower the FLOPs, the faster the model, i.e. the better (less operations). At this point, here, we also note that FLOPs are different from Floating-point operations Per Second (FPS). To be machine GPU agnostic, we use FLOPs and we do not measure time (i.e. seconds). In our FLOPs profiling analysis, we also include the models TerraMind [29] and Prithvi-EO-2.0 [23] (see Sec. V-E).
Sentinel-2, L2A data. For the models we develop, implement, train, and evaluate that are based on the EO FM PhilEO [3], we use images of size 128128x128128, and also $10$ spectral bands (channels), i.e. multi-spectral satellite Sentinel-2 data. More specifically, as in [3], we use the $10$ and $20$ metre spatial resolution spectral bands from the satellite Sentinel-2555http://sentiwiki.copernicus.eu/web/s2-mission#S2Mission-SpatialResolutionS2-Mission-Spatial-Resolutiontrue.
@@
Figure 8: Road and building density estimation: Evaluation in RMSE and also average over the different $n$-shots, examining the FLOPs needed by the models, where the best performing model is the PhilEO Geo-Aware U-Net. We also note that the two EO FM models TerraMind [29] and Prithvi-EO-2.0 [23] are also evaluated in Sec. V-E for roads, buildings, and land cover.
Figure 9: Land cover mapping: Evaluation in accuracy and average over the $n$-shots, examining the FLOPs of the models, where the best model is TerraMind.
V-C Comparison of scaling-up for U-Net and ViT models
Road density estimation. Fig. 7 shows the RMSE performance for road density estimation across different $n$-shot settings. For all $n$-shots, the scaled-up PhilEO MajorTOM 23TB model outperforms the Geo-Aware U-Net model pretrained on the PhilEO Globe 0.5TB dataset. At $n=100$, the PhilEO MajorTOM 23TB model achieves a RMSE of $0.06724$, compared to $0.07070$ for the model pretrained on the PhilEO Globe 0.5TB dataset. This corresponds to an improvement of $4.89\%$. In the figure, we denote the Geo-Aware U-Net by U-Net for clarity purposes, where the Geo-Aware U-Net that we train and use is a modified U-Net pretrained on S-2 global data. For $n=50$ and $100$, the PhilEO 44M MajorTOM model outperforms the FastTOM and PhilEO Globe pretrained baselines, demonstrating the benefits of large-scale pretraining even when fine-tuning with limited labeled data.
@@ V-C Comparison of scaling-up for U-Net and ViT models
  • ViT 100M FastTOM: a Vision Transformer model (approximately 100M parameters, with depth $32$, number of heads $16$, and embedding dimension $512$) pretrained on FastTOM with a UPerNet decoder utilizing intermediate layers $5$, $15$, $23$, and $31$.
@@
Evaluation results, and land cover being correlated with seasonality. The dataset PhilEO Globe 0.5TB is multi-temporal and includes four time steps corresponding to the four seasons of the year. In Fig. 11, we observe much more correlation between land cover and including the four seasons of the year in the training set compared to between road density and including the four seasons in the training dataset, as well as between building density and including seasonality in the pretraining dataset, in Figs. 7 and Fig. 10 respectively.
@@
  • Scaling effect: How does increasing the pretraining dataset size from 0.5TB and 2TB (specialized) to 23TB (general-purpose) impact downstream performance?
@@ V-D Comparison of scaling-up for Mamba and ViT models
Evaluation of using larger patch size, and efficiency-performance trade-off. In Fig. 17, as well as in Figs. 15 and 16, we also evaluate, plot, and examine the performance of the model when we use patches of larger size, i.e. of more pixels. Using different patch sizes, we study the trade-off between efficiency (i.e. larger patches, more coarse-grained and faster) and performance. We examine ViT UPerNet P16 100M-2T and ViT UPerNet P16 300M ImageNet. Regarding the latter, we examine the model size Large which has a bigger number of model parameters (300M), i.e. an interesting comparison. In addition to comparisons and discussion about the patch size, we also have comparisons and discussion about the pretraining dataset, i.e. EO-specific or ImageNet, as well as about the model size. The ViT UPerNet P16 100M model has 19.1219.12G FLOPs, while ViT UPerNet P16 300M ImageNet has 48.4848.48G FLOPs. We observe that pretraining on satellite S-2 data (i.e. rather than on natural images, ImageNet) is beneficial for for low $n$-shot samples. For example, this is also true for patches of size $8$, i.e. for ViT UPerNet P8 100M, where for land cover mapping, for the $n$-shot $500$, the model pretrained on S-2 data has an accuracy of $0.5713$, while the model pretrained on ImageNet $0.5343$ (i.e. percentage improvement $6.92\%$).

### v3

@@ I Introduction
Our scaling-up study. In this paper, we present the scaling-up of our recently proposed model PhilEO on the unlabeled 23TB dataset MajorTOM [15], which covers the vast majority of the Earth’s surface, as well as on the specialized subset FastTOM 2TB that does not include oceans and ice. FastTOM is a task-specific dataset for land only. In the literature, there exist FMs only for oceans or ice, for example HydroFM [25], but we are interested in Earth Observation and not only in Land Observation or Ocean Observation (or Ice Observation). To this end, in this work, we develop and study various models with different numbers of parameters, and also with different architectures. We evaluate and examine the FLoating-point OPerations (FLOPs) that the models need [65]. We fine-tune the models on the PhilEO Bench [3] for road density estimation, building density pixel-wise regression, and land cover semantic segmentation, and we evaluate their performance.
@@
  • We train our model on the pretraining dataset MajorTOM 23TB which includes all regions, and the performance is competitive versus models pretrained on more specialized datasets (i.e. substantially smaller datasets that only include land). The additional data of oceans and ice on average do not decrease the performance on land-focused downstream tasks. These results indicate that large GFMs trained on global datasets for a wider variety of downstream tasks can be useful for downstream applications that only require a subset of the information included in their training.
  • The second contribution is the exploration of U-Net, ViT, and Mamba as FMs for EO. U-Net captures local correlations amongst pixels, while ViT and Mamba capture local and distant correlations amongst patches (or pixels).
We scale the pretraining from 0.5TB to 23TB, as well as to 2TB, where the 23TB are general-purpose EO data. We use different number of model parameters and architectures [^1] We release the code of our model in: http://github.com/ESA-PhiLab/PhilEO-MajorTOM.. We show that for road density regression, for the $n$-shot $n=500$, larger pretraining data reduces the final downstream task’s Root Mean Squared Error (RMSE), resulting in a percentage improvement $12.0\%$. In addition to pretraining dataset scaling, we also study the scaling-up of the model parameters and the architecture, comparing CNN- and ViT-based models, as well as Mamba. We evaluate the effectiveness of the scaling-up on the PhilEO Bench, covering both road and building density estimation, as well as land cover mapping.
We also develop and compare different architectures for the EO Foundation Models in order to select which one to scale-up. More specifically, for road density regression, we show that the ViT UPerNet model outperforms the ViT with a convolutional decoder in RMSE, with a percentage improvement of $34.2\%$ for the $n$-shot $n=100$. On the same task, the U-Net CNN Geo-Aware model achieves the same performance as the ViT UPerNet.
@@ II Background
MajorTOM and FastTOM datasets. The MajorTOM dataset [15] comprises in total approximately 60TB of unlabeled global data, covering most of the Earth, including oceans and ice. In this work, we focus on the MajorTOM Core-S2L2A subset, 23TB. In addition, FastTOM is a 2TB specialized subset of MajorTOM Core-S2L2A [^2] http://github.com/ESA-PhiLab/Major-TOM and also in http://huggingface.co/Major-TOM, containing only land and excluding oceans and ice, making it more task-specific for terrestrial downstream tasks.
@@ III-A PhilEO MajorTOM 23TB and PhilEO FastTOM 2TB
In this subsection, we note that our main contribution is pretraining on the FastTOM and MajorTOM datasets. For the pretraining, the models we develop, implement, and train have three output heads [3], one for reconstruction with the Mean Squared Error (MSE) loss where we mask a percentage of the input image during training, one for estimating the geo-location longitude and latitude where cosine and sine values are used, and one for predicting the Köppen-Geiger climate zone classes [3], [67]. “Geo-Aware” (for example, in the PhilEO Geo-Aware U-Net model) refers to these three output heads, i.e. to the reconstruction objective, the geo-location estimation, and the Köppen-Geiger climate zone prediction.
@@ III-B PhilEO ViT UPerNet
ViT UPerNet architectures have been shown to achieve state-of-the-art performance in various computer vision tasks [11]-[14], including segmentation and classification. For models that use a ViT backbone, we extract intermediate feature maps, $\{C_{2},C_{3},C_{4},C_{5}\}$ in Fig. 1, across different layers. These feature maps are then appropriately resized (i.e. upsampled or downsampled) to align spatial dimensions.
The highest-level feature map $C_{5}$ is processed by the PPM, which applies pooling operations at multiple scales to capture contextual information at different resolutions. The outputs of the PPM and the FPN are then fused through a series of 1×1 convolutions and element-wise additions, producing the final set of multi-scale feature maps, $\{P_{2},P_{3},P_{4},P_{5}\}$ in Fig. 1. These aggregated features are then used for downstream pixel-wise prediction tasks, including regression and segmentation. Finally, in this subsection, we note that our main contributions are: i) including the ViT UPerNet in the PhilEO Bench (i.e. see Sec. V-A, as well as Figs. 4-6), and ii) studying the scaling-up of the ViT UPerNet model (see Sec. V-E, as well as Figs. 15-17, and also Sec. V-C, as well as Figs. 7-11).
@@ III-C Mamba 2D Models
ViT models have significantly less inductive bias compared to CNN U-Net models. This is because ViT attention correlates everything with everything else. Moreover, in self-attention, image token order does not matter [9], [30]. In addition, Mamba models have more image-inductive bias compared to ViT models, as in Mamba S6 SSM models, the image token order is important and matters. In addition, we also note that our main contribution in this subsection is including Mamba SSM 2D models in the PhilEO Bench (i.e. see Sec. V-D).
@@ IV Related Work
Recent studies on GFMs, summarized in Table I, highlight the increasing potential of learning robust feature representations from large-scale EO data, later used for fine-tuning on various downstream tasks. EO FMs can be trained, for example, using MAE [1], [56], [57], contrastive learning [2], or Joint Embedding Predictive Architecture (JEPA) [39]. MAE refers to masking (e.g., a large percentage of the image such as $75\%$), using the reconstruction loss, and the MSE [23], [26]. Also, the ViT [9] (or a variant of the ViT like the Swin Transformer [12]) is usually used. Here, we also note that the Spectral Angle Mapper (SAM), which takes into account the spectral signatures of the observations [^4] http://step.esa.int/main/wp-content/help/versions/9.0.0/snap-toolboxes/org.esa.s2tbx.s2tbx.spectral.angle.mapper.ui/sam/SAMProcessor.html, can also be used in addition to the MSE, or even on its own as the loss function that is minimized during training. In addition, contrastive learning refers to performing data augmentation and bringing closer together in the latent feature space the data-augmented samples, repelling the dissimilar samples.
For a recent study on the scope of geo-spatial FMs, we refer the reader to [60]. However, most (but certainly not all) existing studies have been limited by relatively modest dataset sizes and have not explored training GFMs at the scale of more than 20TB. This motivates our work, where we develop and pretrain the PhilEO MajorTOM model on the full 23TB MajorTOM Core-S2L2A dataset. Furthermore, during the ESA-NASA Workshop on AI Foundation Models for Earth Observation [61] (in May 2025), it was highlighted that the performance of FMs is not uniform across tasks; moreover, in certain cases, conventional models surpass state-of-the-art EO FMs. This observation prompts a critical inquiry: should the field pursue a single universal model, develop application-specific models, or adopt a Mixture of Experts (MoE) strategy?
@@ V Evaluation and Results
The experiments in this paper are as follows. We have three datasets, 0.50.5TB, 22TB, and 2323TB, and we also have three types of models, U-Net CNN, ViT UPerNet, and Mamba. In addition, we have three downstream tasks. We first present the 0.50.5TB experiments in Sec. V-A, i.e. Figs. 4-6. Next, we present the U-Net CNN and ViT experiments, i.e. in Figs. 7-11 (Sec. V-C). Here, the U-Net 2323TB scaling-up experiments are presented. We then present the Mamba results (i.e. Figs. 12-14). These are in Sec. V-D. Subsequently, we present the ViT UPerNet 2323TB scaling-up experiments, in Figs. 15-17 (i.e. Sec. V-E). Next, we present the comparison with TerraMind [29] and Prithvi-EO-2.0 [23]. This is in Figs. 18-20. We are doing these experiments in this order because in the beginning, for the 0.50.5TB results, we examine which architecture/ model to scale-up based on the results on the small dataset. Next, for the U-Net CNN versus ViT experiments, this is an interesting comparison because U-Net CNNs capture local correlations while ViTs capture global, including local, correlations. Then, for the ViT versus Mamba experiments, this is an interesting comparison because they both capture global correlations and Mamba is more computationally efficient. For the ViT 2323TB scaling-up experiments, this is interesting because we increase the pretraining dataset size and study what happens when oceans and ice data are included in the pretraining. Also, for the comparison with TerraMind and Prithvi experiments, this is interesting because they are recent state-of-the-art models. The main findings (as we will also see in the sections below) are that the Geo-Aware U-Net 200200M model is effective for pixel-wise regression tasks and outperforms the other models. In addition, the FLOPs analysis indicates that the smaller model U-Net 4444M can be effective for pixel-wise regression downstream tasks. Furthermore, the comparison with TerraMind shows that for pixel-wise regression tasks, for some $n$-shot samples, the proposed models can outperform TerraMind.
Figure 4: Road density estimation: Evaluation in RMSE at different $n$-shots, where the best performing model is PhilEO Geo-Aware U-Net. The PhilEO ViT UPerNet black line is below the Geo-Aware U-Net red line. ViT UPerNet outperforms ViT CNN and ViT CNN Grouped Channels (GC) [3], [7].
@@ V-B Scaling-up: Evaluation of PhilEO MajorTOM 23TB GFM
In this section, we present the evaluation and results of the models when we scale-up FMs in terms of larger pretraining dataset (MajorTOM Core-S2L2A 23TB dataset), larger model, and architecture. According to the results, evaluation, and experiments, FMs pretrained on general-purpose data, when scaled-up appropriately, can eventually outperform models pretrained on specialized data. We are interested in Earth Observation, and not only in Land Observation or Ocean Observation (or Ice Observation). In the literature, there exist FMs only for oceans or ice, for example HydroFM [25]. When we pretrain on the entire MajorTOM 23TB dataset, we do not use any prior information, while when we pretrain on specialized datasets that have only land and exclude oceans, ice, and deserts, i.e. the FastTOM 2TB dataset, then we use prior information. The downstream tasks we are interested in are on land, i.e. the PhilEO Bench [3]: road and building density estimation, and land cover mapping [^5] The PhilEO Bench downstream tasks labeled dataset (0.40.4TB) was presented in [3], and we also note that a small and more manageable subset can be found in http://huggingface.co/datasets/JenteBosmans/tiny_phileo.. We focus on land downstream tasks [^6] http://phileo-bench.github.io/ and not on oceans or ice (or deserts), and this is why specialized datasets have prior knowledge.
@@
For the evaluation of the models, continuing the discussion in the preceding paragraph, we count the FLOPs that are needed by the different models [65]. The lower the FLOPs, the faster the model, i.e. the better (less operations). We measure the FLOPs in Python using the DeepSpeed library. Here, we also note that FLOPs are different from Floating-point operations Per Second (FPS). To be machine GPU agnostic, we use FLOPs and we do not measure time (i.e. seconds). In our FLOPs profiling analysis, we also include the models TerraMind [29] and Prithvi-EO-2.0 [23] (see Sec. V-E).
Figure 8: Road and building density estimation: Evaluation in RMSE and also average over the different $n$-shots, examining the FLOPs needed by the models [65], where the best performing model is the PhilEO Geo-Aware U-Net. We also note that the two EO FM models TerraMind [29] and Prithvi-EO-2.0 [23] are also evaluated in Sec. V-E for roads, buildings, and land cover.
Figure 9: Land cover mapping: Evaluation in accuracy and average over the $n$-shots, examining the FLOPs of the models [65], where the best is TerraMind.
Sentinel-2, L2A data. For the models we develop, implement, train, and evaluate that are based on the EO FM PhilEO [3], we use images of size 128128x128128, and also $10$ spectral bands (channels), i.e. multi-spectral satellite Sentinel-2 data. More specifically, as in [3], we use the $10$ and $20$ metre spatial resolution spectral bands from the satellite Sentinel-2777http://sentiwiki.copernicus.eu/web/s2-mission#S2Mission-SpatialResolutionS2-Mission-Spatial-Resolutiontrue.
@@ V-C Comparison of scaling-up for U-Net and ViT models
Results for road density estimation. Fig. 7 shows the RMSE for road density estimation across different $n$-shot settings. For all $n$-shots, the scaled-up PhilEO MajorTOM 23TB model outperforms the Geo-Aware U-Net model pretrained on the PhilEO Globe 0.5TB dataset. At $n=100$, the PhilEO MajorTOM 23TB model achieves a RMSE of $0.06724$, compared to $0.07070$ for the model pretrained on the PhilEO Globe 0.5TB dataset. This corresponds to an improvement of $4.89\%$. In the figure, we denote the Geo-Aware U-Net by U-Net for clarity purposes, where the Geo-Aware U-Net that we train and use is a modified U-Net pretrained on S-2 global data. For $n=50$ and $100$, the PhilEO 44M MajorTOM model outperforms the FastTOM and PhilEO Globe pretrained baselines, demonstrating the benefits of large-scale pretraining even when fine-tuning with limited labeled data.
@@
  • ViT 100M FastTOM: a Vision Transformer model (approximately 100M parameters, with depth $32$, number of heads $16$, and embedding dimension $512$) pretrained on FastTOM with a UPerNet decoder utilizing the intermediate layers $5$, $15$, $23$, and $31$.
@@
Evaluation results, and land cover being correlated with seasonality. The dataset PhilEO Globe 0.5TB is multi-temporal and includes four time steps corresponding to the four seasons of the year. In Fig. 11, we observe much more correlation between land cover and including the four seasons of the year in the training set, compared to between road density and including the four seasons in the training dataset, as well as between building density and including seasonality in the pretraining dataset, in Figs. 7 and Fig. 10 respectively.
@@
  • Scaling effect: How does increasing the pretraining dataset size from 0.50.5TB and 22TB (specialized) to 2323TB (general-purpose) impact downstream performance?
@@ V-D Comparison of scaling-up for Mamba and ViT models
Evaluation of using larger patch size, and efficiency-performance trade-off. In Fig. 17, as well as in Figs. 15 and 16, we also evaluate, plot, and examine the performance of the model when we use patches of larger size, i.e. of more pixels. Using different patch sizes, we study the trade-off between efficiency (i.e. larger patches, more coarse-grained and faster) and performance. We examine ViT UPerNet P16 100M-2T and ViT UPerNet P16 300M ImageNet. Regarding the latter, we examine the model size Large which has a bigger number of model parameters (300M), i.e. an interesting comparison. In addition to comparisons and discussion about the patch size, we also have comparisons and discussion about the pretraining dataset, i.e. EO-specific or ImageNet, as well as about the model size. The ViT UPerNet P16 100M model has 19.1219.12G FLOPs, while ViT UPerNet P16 300M ImageNet has 48.4848.48G FLOPs. We observe that pretraining on satellite S-2 data (i.e. rather than on natural images, ImageNet) is beneficial for low $n$-shot samples. For example, this is also true for patches of size $8$, i.e. for ViT UPerNet P8 100M, where for land cover mapping, for the $n$-shot $500$, the model pretrained on S-2 data has an accuracy of $0.5713$, while the model pretrained on ImageNet $0.5343$ (i.e. percentage improvement $6.92\%$).